Grade 12th passPhysical Chemistry

Which has the maximum number of moleculesa)64 g of SO2b)44 g of CO2c)48 g of O2d)8 g of H2

Omkar Jangale
8 Years agoGrade 12th pass

1 Answer

Rishabh Gaikwad
8 Years ago
a) Molar Mass SO2 = 32 + 2(16)= 64g
   Therefore, mole = Given mass / Molar mass 
   = 64/64 = 1 mole
   Therefore, molecule = mole * NA
   = 1 * 6.022 x 1023
    = 6.022 x1023 molecules.
b) Molar mass of Co2 = 12 + 2(16) = 44g
 Therefore, mole = Given mass / Molar mass
 = 44/44 = 1 mol
   Therefore, molecule = mole * NA
= 1 * 6.022 x 1023  
= 6.022 x1023 molecules.
c) Molar mass of O2 = 2(16) = 32g
Therefore, mole = Given mass / Molar mass 
= 48/32 = 3/2 mole
=Therefore, molecule = mole * NA
= 3/2 * 6.022 x 1023  
=9.033 x1023 molecules.
​d) Molar mass of H2 = 2(1) = 2g
Therefore, mole = Given mass / Molar mass 
= 8/2= 4 mole
=Therefore, molecule = mole * NA
= 4 * 6.022 x 1023  
=24.088 x1023 molecules.